### v3

# Aligning Text to Image in Diffusion Models is Easier Than You Think

Abstract
While recent advancements in generative modeling have significantly improved text-image alignment, some residual misalignment between text and image representations still remains. Although many approaches have attempted to address this issue by fine-tuning models using various reward models, etc., we revisit the challenge from the perspective of representation alignment—an approach that has gained popularity with the success of REPresentation Alignment (REPA) [36]. We first argue that conventional text-to-image (T2I) diffusion models, typically trained on paired image and text data (i.e., positive pairs) by minimizing score matching or flow matching losses, is suboptimal from the standpoint of representation alignment. Instead, a better alignment can be achieved through contrastive learning that leverages both positive and negative pairs. To achieve this efficiently even with pretrained models, we introduce a lightweight contrastive fine-tuning strategy called SoftREPA that uses soft text tokens. This approach improves alignment with minimal computational overhead by adding fewer than 1M trainable parameters to the pretrained model. Our theoretical analysis demonstrates that our method explicitly increases the mutual information between text and image representations, leading to enhanced semantic consistency. Experimental results across text-to-image generation and text-guided image editing tasks validate the effectiveness of our approach in improving the semantic consistency of T2I generative models. Project Page: https://softrepa.github.io/
Figure 1: Representative results for image generation and image editing. SoftREPA provides much improved text to image alignment by introducing a negligible size of learnable soft tokens.
1 Introduction
Achieving effective alignment between modalities is essential in multimodal generative modeling. Specifically, latent diffusion models [26] enable various conditioning mechanisms, such as text, during the generation process for the semantic alignment between the two modalities.
For UNet-based denoisers [26, 22, 24], image representations are updated to align with fixed text representations encoded by pre-trained CLIP models through cross-attention layers. In contrast, transformer-based denoisers [23, 7] jointly update both text and image representations by concatenating them and processing them through self-attention layers. However, there is still room for further improvement, and to mitigate the remaining misalignment between text and image representations, a more effective representation learning approach is necessary.
Recent advancements in representation alignment for Diffusion Transformers (DiT) [23] have significantly enhanced their ability to learn semantically meaningful internal representations [36, 9]. Notably, REPA [36] demonstrated that aligning the internal representations of DiT with an external pre-trained visual encoder during training significantly improves both discriminative and generative performance.
Figure 2: Network architecture and algorithmic concept of SoftREPA. (a) Learnable soft tokens are prepended to the text features across the upper layers, with each layer’s previous soft tokens being replaced by the soft tokens of the current layer. (b) The soft tokens are optimized to contrastively match the score with positively conditioned predicted noise while repelling the score from negatively conditioned predicted noise. This process implicitly sharpens the joint probability distribution of images and text by reducing the log probability of negatively paired conditions.
To enhance the textual semantic alignment of DiT’s internal representations, we leverage these ideas to further enhance text-image alignment in text-to-image (T2I) generative models. Specifically, we are interested in the contrastive learning framework, a widely used strategy for multimodal alignment in most vision-language representation learning approaches such as CLIP [25], ALBEF [15], BLIP [16], ALIGN [10].
Most vision-language generative foundation models have focused on improving multimodal alignment through architectural modifications. These include MLP-based projection layers (e.g., LLaVA-style visual instruction tuning) [19, 20], cross-attention-based fusion mechanisms (e.g., Flamingo, Stable Diffusion 1.5) [2, 22], and early fusion models that integrate text and vision features within a shared representation space (e.g., Chameleon, Stable Diffusion 3, FLUX) [30, 7]. In fact, many of these approaches have been largely adopted in T2I generative models. However, we explore an efficient yet effective way of further enhancing the representation alignment for given pre-trained models.
One of the most important contributions of this work is therefore the introduction of soft text tokens which are trained through contrastive text-and-image training. This allows the model to dynamically adapt its text representations, improving alignment with generated images without requiring full model finetuning. This simple yet effective approach significantly enhances text-to-image alignment in text-to-image generation and text-guided image editing tasks. Furthermore, the method is so flexible that it can be used with any pretrained T2I generative models. Theoretically, we show that this approach explicitly increases the mutual information between text and images, leading to better semantic consistency in multi-modal representations. We therefore refer to our method as SoftREPA, short for Soft REPresentation Alignment. Our contributions can be summarized as follows.
  • We propose SoftREPA, a novel text-image representation alignment method that leverages a lightweight fine-tuning strategy with soft text tokens. This approach improves text-image alignment while adding fewer than 1M additional parameters, ensuring efficiency with minimal computational overhead.
  • SoftREPA is simple yet flexible so that it can be used with any pretrained T2I generative models to improve performance of image generation, editing, etc.
  • We show that our method explicitly increases mutual information between image and text, leading to better semantic consistency in multi-modal representations.
2 Preliminaries
Flow Models. Suppose that we have access to samples from target distribution $X_{0}\sim q$ and source distribution $X_{1}\sim p$. The goal of flow model is to generate $X_{0}$ starting from $X_{1}$. Specifically, we define a velocity field $v_{t}({\bm{x}})$ of a flow $\psi_{t}({\bm{x}}):[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that satisfies $\psi_{t}(X_{0})=X_{t}$ and $\psi_{1}(X_{0})=X_{1}$. Here, the $\psi_{t}$ is uniquely characterized by a flow ODE:
$d\psi_{t}({\bm{x}})=v_{t}(\psi_{t}({\bm{x}}))dt$ (1)
where the flow velocity $v_{t}$ is fitted to the parameterized neural network $v_{t,\theta}$ via flow matching:
$\mathcal{L}_{FM}=\mathbb{E}_{t\in[0,1],{\bm{x}}_{t}\sim p_{t}}\|v_{t}({\bm{x}}_{t})-v_{t,\theta}({\bm{x}}_{t})\|^{2}.$ (2)
However, this is computationally expensive to solve due to the integration with respect to $X_{0}$, so the authors in [18] proposed a conditional flow matching that has the same gradient with the original objective function:
$\mathcal{L}_{CFM}=\mathbb{E}_{t\in[0,1],{\bm{x}}_{0}\sim q}\|v_{t}({\bm{x}}_{t}|{\bm{x}}_{0})-v_{t,\theta}({\bm{x}}_{t})\|^{2},$ (3)
where $v_{t}({\bm{x}}_{t}|{\bm{x}}_{0})$ is a conditional velocity field that defines a conditional flow $\psi_{t}({\bm{x}}_{t}|{\bm{x}}_{0})$ satisfying $\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})={\bm{x}}_{t}$.
In particular, the linear conditional flow defines the flow as ${\bm{x}}_{t}=\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})=(1-t){\bm{x}}_{0}+t{\bm{x}}_{1}$. Then, we can compute the conditional velocity field $v_{t}({\bm{x}}_{t}|{\bm{x}}_{0})=\dot{\psi}_{t}(\psi_{t}^{-1}({\bm{x}}_{t}|{\bm{x}}_{0})|{\bm{x}}_{0})={\bm{x}}_{1}-{\bm{x}}_{0}$ [18]. Thus, the conditional flow matching loss is defined as
$\mathbb{E}_{t\in[0,1],{\bm{x}}_{0},{\bm{x}}_{1}\sim\pi_{0,1}}\|({\bm{x}}_{1}-{\bm{x}}_{0})-v_{t,\theta}({\bm{x}}_{t})\|^{2}.$ (4)
Considering a marginal velocity field,
[…]
$=\mathbb{E}[{\bm{x}}_{1}-{\bm{x}}_{0}|{\bm{x}}_{t}]=\mathbb{E}[{\bm{x}}_{1}|{\bm{x}}_{t}]-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}],$
the generation process via flow ODE is
$d{\bm{x}}_{t}=v_{t}({\bm{x}}_{t})dt=(\mathbb{E}[{\bm{x}}_{1}|{\bm{x}}_{t}]-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])dt.$ (6)
Accordingly,
$d{\bm{x}}_{t}$ $=\left(\mathbb{E}\left[\frac{{\bm{x}}_{t}-(1-t){\bm{x}}_{0}}{t}|{\bm{x}}_{t}\right]-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]\right)dt$ (7)
$=\frac{{\bm{x}}_{t}-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]}{t}dt$
which is the probability-flow ODE (PF-ODE) of DDIM [28, 11] when ${\bm{x}}_{1}\sim p:=\mathcal{N}(0,{\mathbf{I}}_{d})$. In other words, flow models and score-based models can be used interchangeably.
2.1 Contrastive Representation Learning
Contrastive learning aims to maximize the similarity between semantically related text-image pairs while pushing apart unrelated pairs in a shared representation space [5, 25]. Formally, given a batch of N image-text pairs ${(\textbf{I}_{i},\textbf{T}_{i})}^{N}_{i=1}$, contrastive learning optimizes a contrastive loss function, typically a variant of the InfoNCE loss:
$\mathcal{L}_{\text{CLIP}}=-\frac{1}{N}\sum^{N}_{i=1}\log\frac{\exp(\text{sim}(\textbf{I}_{i},\textbf{T}_{i})/\tau)}{\sum^{N}_{j=1}\exp(\text{sim}(\textbf{I}_{i},\textbf{T}_{j})/\tau)}$ (8)
where $\text{sim}(\textbf{I}_{i},\textbf{T}_{i})$ is a similarity metric such as cosine similarity or inner product between image and text embeddings, and $\tau$ denotes a temperature parameter that controls the distribution sharpness. This contrastive loss encourages image and text representations to form a joint multimodal representation space, where aligned pairs are close together, and unaligned pairs are separated. Several fundamental models have successfully applied contrastive learning for multi-modal representation learning [25, 10, 15, 16] and self-supervised learning [5].
Despite the success of contrastive learning in representation learning, its application in generative models remains underexplored due to key challenges. First, there is a fundamental mismatch between discriminative and generative features; conventional contrastive learning optimizes a representation space for discriminative tasks, whereas generative models focus on realistic sample synthesis. Second, text-image alignment in diffusion-based generative models is often implicit, relying on denoising objectives rather than explicit representation learning.
To bridge this gap, we propose a contrastive learning framework that aligns both representation learning and generative objectives, ensuring effective text-image representation alignment while preserving generative quality.
3 SoftREPA
3.1 Contrastive T2I Alignment Loss
Let $\{({\bm{x}}^{(i)},{\bm{y}}^{(i)})\}_{i=1}^{n}$ denotes the matched image and text pairs as a training data set. Then, contrastive text-to-image alignment in T2I model can be generally formulated as follows:
$\mathcal{L}=-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\exp(l({\bm{x}}^{(i)},{\bm{y}}^{(i)}))}{\sum_{j}\exp(l({\bm{x}}^{(i)},{\bm{y}}^{(j)}))}$ (9)
where $l(\cdot,\cdot)$ represents the similarity measure. It is important to note that unlike the standard T2I model training that only consider positive pairs, i.e. $({\bm{x}}^{(i)},{\bm{y}}^{(i)})$, our SoftREPA employs the contrastive T2I alignment that additionally considers negative image and text pairs $({\bm{x}}^{(i)},{\bm{y}}^{(j)})),i\neq j$. This significantly improves the alignment performance.
Furthermore, in SoftREPA we define $l({\bm{x}}^{(i)},{\bm{y}}^{(j)})$ as the logit from the variation of the denoising score matching loss for the case of T2I diffusion models [31].
$l({\bm{x}},{\bm{y}})=e^{-\mathbb{E}_{t,{\bm{\epsilon}}}[\left\lVert{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\right\rVert^{2}/\tau(t)]}$ (10)
Similarly, in conditional flow matching, the logit $l({\bm{x}},{\bm{y}})$ can be formalized as:
$l({\bm{x}},{\bm{y}})=e^{-\mathbb{E}_{t,\epsilon}[\|v_{\theta}({\bm{x}}_{t},t,{\bm{y}})-({\bm{\epsilon}}-{\bm{x}}_{0})\|^{2}/\tau(t)]}$ (11)
Here, $\tau(t)$ represents time scheduling parameter. One might consider using the denoising score matching loss, $-\mathbb{E}_{t,{\bm{\epsilon}}}[\left\lVert{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\right\rVert^{2}]$, as a similarity measure instead of relying on its logit value. However, our experiments revealed that this formulation can introduce instability during training due to the unbounded nature of the similarity measure. To address this, SoftREPA introduce an exponential function to constrain the logit values, effectively stabilizing the training process.
Intuitively, as illustrated in Fig. 2(b), training a diffusion model with this loss function would lead to predicted noise with matching text description align with the true noise. Conversely, when conditioned on a mismatched text, the predicted noise would deviate further from the true noise. It implies that the conditional probability distribution of the trained model gets sharpened, increasing the distinction between different text conditions, which is expected to enhance image-text alignment in the generated images.
3.2 The Soft Tokens
To distill the contrastive score matching objective above, we introduce a learnable soft token ${\bm{s}}_{\psi}$, which is the only trainable parameter while pretrained model remains frozen. In particular, as illustrated in Fig. 2, we introduce a set of learnable soft tokens that vary between layers and are indexed by time $t$. The soft token at layer $k$ is defined using an embedding function with the number of soft tokens as $m$:
$\mathbf{s}^{(k,t)}_{\psi}=\text{Embedding($k,t$)}\in\mathbb{R}^{m\times d}$ (12)
At the layer $k$, the text representation is updated by concatenating the time-indexed soft token with text hidden representation from the previous layer: $\hat{\mathbf{H}}^{(k-1,t)}_{\text{text}}=[\mathbf{s}^{(k,t)}_{\psi};\mathbf{H}^{(k-1,t)}_{\text{text}}]\in\mathbb{R}^{(m+n)\times d}$.
The soft token is prepended on text features across the layers, leading to a modified denoising function: $v_{\theta}({\bm{x}}_{t},t,{\bm{y}},{\bm{s}}_{\psi})$. By adopting the soft tokens, we can adjust the image and text representations to get better semantic alignment. Additionally, for computational efficiency, we approximate the expectation within $l({\bm{x}},{\bm{y}})$ using a single Monte Carlo sample, effectively replacing it as
$\tilde{l}({\bm{x}},{\bm{y}},{\bm{s}}_{\psi})=e^{-\|v_{\theta}({\bm{x}}_{t},t,{\bm{y}},{\bm{s}}_{\psi})-({\bm{\epsilon}}-{\bm{x}}_{0})\|^{2}/\tau(t)}$ (13)
which can stabilize the computation using the same ${\bm{\epsilon}}$ and $t$ in the same batch. The resulting contrastive loss function with soft tokens can be defined as follows:
$\mathcal{L}_{\text{SoftREPA}}(\psi)=-\mathbb{E}_{\begin{subarray}{c}({\bm{x}},{\bm{y}})\sim p_{\text{data}}\\ t\sim U(0,1)\\ {\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\end{subarray}}\log\frac{\exp(\tilde{l}({\bm{x}},{\bm{y}},{\bm{s}}_{\psi}))}{\sum_{j}\exp(\tilde{l}({\bm{x}},{\bm{y}}^{(j)},{\bm{s}}_{\psi}))}$ (14)
Then the learnable token ${\bm{s}}_{\psi}$ is optimized as follows:
$\mathbf{s}^{*}_{\psi}=\arg\min_{\psi}\mathcal{L}_{\text{SoftREPA}}(\psi)$ (15)
During inference, the trained soft tokens are leveraged with fixed text representation from the text encoder across layers and timesteps. The detailed procedure during the image generation process based on MM-DiT architecture is described in Algorithm 1.
Algorithm 1 Image Generation with Soft Tokens in MM-DiT
1:Gaussian noise $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
2:Text $\mathbf{y}\sim p_{\text{data}}$
3:Soft token $\mathbf{s_{\psi}}\sim\text{Embedding}(k,t)$
4:Number of layers $N$, Threshold layer $L$
5:Time steps $\{t_{T},t_{T-1},...,t_{0}\}$
6:Initialize $\mathbf{H}_{\text{img}}^{(0,T)}\leftarrow\mathbf{z}$
7:Initialize $\mathbf{H}_{\text{text}}^{(0,T)}\leftarrow\text{TextEncoder}(\mathbf{y})$
8:$n=|\mathbf{H}_{\text{text}}^{(0,T)}|$
9:for $t$ in $\{t_{T},t_{T-1},...,t_{0}\}$ do
10:    for $l=1$ to $N$ do
11:         if $k\leq L$ then
12:             $\mathbf{s_{\psi}}^{(k,t)}\leftarrow\text{Embedding}(k,t)$
13:             $\hat{\mathbf{H}}_{\text{text}}^{(k-1,t)}\leftarrow[\mathbf{s_{\psi}}^{(k,t)};\mathbf{H}_{\text{text}}^{(k-1,t)}]$
14:         else
15:             $\hat{\mathbf{H}}_{\text{text}}^{(k-1,t)}\leftarrow\mathbf{H}_{\text{text}}^{(k-1,t)}$
16:         end if
17:         $\mathbf{H}_{\text{img}}^{(k,t)},\hat{\mathbf{H}}_{\text{text}}^{(k,t)}\leftarrow\text{Layer}_{l}(\mathbf{H}_{\text{img}}^{(k-1,t)},\hat{\mathbf{H}}_{\text{text}}^{(k-1,t)})$
18:         $\mathbf{H}_{\text{text}}^{(k,t)}\leftarrow\hat{\mathbf{H}}_{\text{text}}^{(k,t)}[-n:,:]$triangleright Drop soft tokens
19:    end for
20:end for
21:return $\mathbf{\hat{X}}=\text{Decoder}(\mathbf{H}_{\text{img}}^{(N,t_{0})})$
3.3 Relationship with Mutual Information
Although the relationship between contrastive loss and mutual information has been previously studied, we revisit it here to provide a more complete theoretical perspective. Additionally, we emphasize the critical role of the logit formulation in the score matching loss, highlighting its importance in ensuring stable and effective optimization.
According to [13], pointwise mutual information (PMI) over image ${\bm{x}}$ and text ${\bm{y}}$ pairs can be defined as follows:
$i({\bm{x}},{\bm{y}})=\log\frac{p_{\theta}({\bm{x}}|{\bm{y}})}{p_{\theta}({\bm{x}})}=\log\frac{p_{\theta}({\bm{x}}|{\bm{y}})}{\mathbb{E}_{p({\bm{c}})}[p_{\theta}({\bm{x}}|{\bm{c}})]}$ (16)
where the second equality follows from the definition of conditional expectation. Song et al. [29] and Kong et al. [13] further show that assuming the diffusion model is an optimal denoiser, the conditional likelihood $p_{\theta}({\bm{x}}|{\bm{y}})$ can be described as (see Appendix A):
$p_{\theta}({\bm{x}}|{\bm{y}})=\exp(\hat{l}({\bm{x}},{\bm{y}})),$ (17)
where
[…]
By approximating the denominator of PMI via Monte Carlo sampling, we arrive at:
$i({\bm{x}},{\bm{y}})\approx\log\frac{\exp(\hat{l}({\bm{x}},{\bm{y}}))}{\frac{1}{N}\sum_{\bm{c}}\exp(\hat{l}({\bm{x}},{\bm{c}}))}$ (19)
This formulation closely resembles a cross-entropy loss, where the conditional log-likelihood $\hat{l}({\bm{x}},{\bm{y}})$ can be interpreted as logits.
From the fact that mutual information is the expectation of the pointwise mutual information, we derive:
$I(X,Y)=\frac{1}{n}\sum_{i=1}^{n}\log\frac{\exp(\hat{l}({\bm{x}}^{(i)},{\bm{y}}^{(i)}))}{\sum_{j}\exp(\hat{l}({\bm{x}}^{(i)},{\bm{y}}^{(j)}))}+D$ (20)
where ${\bm{x}}^{(i)}$ represents the $i$-th sample in the data, $n$ denotes the number of samples in the data, and $D$ indicates the constant.
Notably, Eq. (20) is very similar in form to the contrastive objective in Eq. (9), where the logit is the negative of the diffusion loss. This implies that minimizing our contrastive learning objective is closely related to maximizing the mutual information between image-text pairs under the diffusion model.
4 Experiments
Implementation details. In our experiments, to confirm the flexibility of SoftREPA, we utilize various open-source text-to-image diffusion models, including Stable Diffusion 1.5, Stable Diffusion XL, and Stable Diffusion 3 to evaluate text-to-image alignment in both image generation and editing tasks. The batch size was set to 16, and we trained the model for fewer than 30,000 iterations using two A100 GPUs.
For Stable Diffusion 3, we set the length of the soft token to 4 and used 5 layers to attach the soft tokens for further experiments. In the case of Stable Diffusion 1.5 and Stable Diffusion XL, the soft token length was set to 8, and soft tokens were applied only to the Down and Middle block layers of the UNet’s original embedding. Refer to further implementation detail in Appendix B.
Figure 3: The qualitative results of text-to-image generation comparing SD3 and SD3 with proposed method. The given text is from COCO and Pixart dataset.
Text to Image Generation
Human Preference Text Alignment Image Quality
[…]
SD1.5 0.177 0.265 0.251 0.438
SD1.5 (Ours) 0.274 0.271 0.252 0.437
SDXL 0.750 0.267 0.273 0.420
SDXL (Ours) 0.852 0.268 0.283 0.423
SD3 0.942 0.263 0.280 0.424
SD3 (Ours) 1.085 0.269 0.289 0.428
Table 1: Quantitative results of image generation regarding image quality and text alignment of generated images.
We conducted text-to-image generation experiments on SD1.5, SDXL, and SD3, training soft tokens using the text-image paired COCO dataset [17]. The implementation details are provided in Appendix B. To evaluate the generated images, we assessed human preference scores [35, 12], text-image alignment [38, 34], and image quality [37, 27] on COCO-val 5K dataset. As shown in Tab. 1, the proposed method generally outperforms baseline approaches in both diffusion and rectified flow model regardless of the model architecture. In particular, detailed textual descriptions are more accurately reflected in the generated images, as demonstrated in the qualitative results (Fig. 3).
Figure 4: The qualitative results of text guided image editing comparing SD3 and SD3 with proposed method. The FlowEdit [14] is used as the editing method for both baseline and SoftREPA.
Figure 5: Quantitative comparisons between baseline editing and editing with soft tokens on DIV2K and Cat2Dog dataset. We evaluate CLIP score and LPIPS for each method. Lines connected with dots represent different CFG scales of the same method.
Human Preference Text Alignment Image Quality
[…]
[14] DIV2K 0.380 0.260 0.255 0.154
[14] (Ours) DIV2K 0.466 0.263 0.260 0.149
[14] Cat2Dog 0.937 0.225 0.266 0.199
[14] (Ours) Cat2Dog 1.111 0.226 0.269 0.173
Table 2: Quantitative evaluation of image editing performance in terms of image quality and target text alignment. The editing process follows FlowEdit [14], with the only difference being the use of soft tokens.
Figure 6: Additional qualitative results of text guided image editing on Cat2Dog dataset.
Text Guided Image Editing
To further demonstrate the improved text-image alignment achieved by our method, we conducted text-guided image editing experiments. These experiments are based on FlowEdit [14], an image editing framework for flow-matching models built on SD3. For the image editing dataset, we selected 800 high-quality images from the DIV2K dataset [1] and generated detailed captions for the source and target images using LLaVA [19]. The dataset generation process is described in detail in  Appendix C.
According to FlowEdit [14], the optimal CFG scale for the target text condition is set to 13.5, with 33 editing steps out of 50 total timesteps. In contrast, with our proposed soft tokens, the CFG scale is reduced to 11, and the number of editing steps is reduced to 30, suggesting that our method could provide a more efficient way to capture the joint text-image distribution. In Fig. 4, the edits in both single and multi-concept scenarios are more effectively reflected in the modified images compared to vanilla SD3. Additional qualitative results are conducted on Cat2Dog dataset Fig. 6 with target CFG scale as 7.
Furthermore, the incorporation of soft tokens leads to improvements in both text-image alignment and overall image quality (Fig. 5). The quantitative results of human preference scores, text alignment, and perceptual image quality evaluation show consistent enhancements with the use of soft tokens (Tab. 2). Detailed results for various parameter settings are provided in Appendix E.
Figure 7: Ablation study analyzing the effect of the number of layers incorporating soft text tokens and the length of tokens. The evaluation is conducted using ImageReward, CLIP score and LPIPS. Each metric is normalized within its respective range: ImageReward $[0.7,1.1]$, CLIP score $[0.25,0.28]$, and LPIPS $[0.42,0.44]$.
Ablation Study
We conducted an ablation study to assess the impact of both the number of layers incorporating soft text tokens and the length of these tokens on the text-to-image generation quality of SD3 using COCO val dataset [17] (Fig. 7). A value of 0 corresponds to the vanilla SD3 model, and the number of layers is counted from the first layer. Each comparison is conducted with either a fixed token length of 8 or a fixed number of 5 layers. Our findings indicate that incorporating soft tokens beyond layer $7$ severely degrades image generation quality. The optimal performance, aligned with human preferences (ImageReward [35]), was observed when soft tokens were applied within layers $2\sim 5$. This aligns with the model’s architectural design, where early layers primarily facilitate text-image alignment, while later layers focus on enhancing image fidelity [21]. However, text-image alignment scores (CLIP) continued to improve as more layers were tuned with soft tokens, which means that the model strongly follows the text conditions as more layers adopt the soft tokens while overall image quality remained stable across different layer configurations. The supporting qualitative results are presented in Appendix F.
Regarding token length, we found that using an excessive number of tokens, particularly beyond a length of 8, negatively impacted image generation quality. This suggests that longer token representations may lead to overfitting on the fine-tuning dataset (Fig. 7). Shorter token lengths ($1\sim 4$ tokens) were beneficial for preserving perceptual similarity to the source dataset (LPIPS), whereas moderate token lengths ($4\sim 8$ tokens) significantly enhanced text-image alignment, as indicated by CLIP scores and ImageReward. Qualitative results supporting these findings are presented in Appendix F.
5 Related Works
Several methods have been proposed to address the text-image misalignment. Training free approach such as CFG++ [6] addresses this problem as the off-manifold phenomenon of conventional CFG and reformulates it to a manifold-constrained CFG to get better text-to-image alignment. Other training-free approach, attend-and-excite [4], adjust attention activations to enhance text-image consistency. DOODL [32] applies guidance to end-to-end manner. Meanwhile, training-based approaches such as Diffusion-DPOK [33] introduce alignment losses during training to both enforce stronger text-conditioning and visual realism. Following Diffusion-DPO, denoising diffusion policy optimization(DDPO) [3] enhances the alignment using feedback from the pre-trained vision-language model. However, these methods require either extensive retraining or an additional human annotation dataset. In this paper, we use the prior knowledge of diffusion model itself to get better text and image alignment, proposing a novel SoftREPA on soft tokens.
6 Conclusion
In this work, we introduce SoftREPA, a novel approach for improving text-image alignment in generative models. Building on the success of prior representation alignment techniques, we extend the idea of internal representation alignment to the text-image alignment without relying on external encoders. Our method effectively integrates contrastive learning and score matching loss to enhance text-image representation learning while preserving the generative capabilities of diffusion models. Also, we introduced soft text tokens, which can effectively adjust the image-text representations through contrastive T2I alignment, making it lightweight and efficient. We demonstrate that our approach enhances text-image alignment in both image generation and editing tasks while maintaining a minimal parameter overhead. Our experiments span various models, including Stable Diffusion 1.5, XL, and 3.
Supplementary Materials
Appendix A Information Theory on Diffusion Model
Several studies have explored diffusion models from an information-theoretic viewpoint. According to [13], the mutual information between random variables $X$ and $Y$ can be expressed as the expectation of pointwise mutual information over their joint distribution:
$I(X,Y)=\mathbb{E}_{p({\bm{x}},{\bm{y}})}[i({\bm{x}},{\bm{y}})],$ (21)
where the pointwise mutual information (PMI) is defined as the difference between the conditional and marginal log-likelihoods of a given sample ${\bm{x}}$:
$i({\bm{x}},{\bm{y}})=\log p({\bm{x}}|{\bm{y}})-\log p({\bm{x}}).$ (22)
This quantity measures how much the presence of condition ${\bm{y}}$ affects the probability distribution near ${\bm{x}}$.
Song et al. [29] and Kong et al. [13] showed that, under the assumption of an optimal diffusion model trained on given data, the log-likelihood of an image ${\bm{x}}$ can be formulated as:
[…]
This result also extends to conditional diffusion models, where the likelihood of ${\bm{x}}$ given condition ${\bm{y}}$ (e.g., text embeddings) is given by:
[…]
Appendix B Implementation Details
The training implementation details are shown in  Tab. 3. For Stable diffusion 3, the length of soft token as $4$ and the number of layers as $5$ are used for the experiments. For Stable diffusion 1.5 and Stable diffusion XL, we used the length of soft token as $8$. Soft tokens are added to the original embedding only on the Down and Middle block layers of UNet. Since we found that the performance of the Stable Diffusion 1.5 degrades when soft tokens are initialized with a random distribution. Therefore, unconditional text embedding was utilized for initializing soft tokens.
In Stable diffusion 3, we optimize the soft tokens only using the contrastive score matching loss ($L_{\text{SoftREPA}}$). In Stable diffusion 1.5 and Stable diffusion XL, we optimize the soft tokens with the combination of the contrastive score matching loss function ($L_{\text{SoftREPA}}$) and the small weight of denoising score matching loss. The code is available at https://github.com/softrepa/SoftREPA.
Models lr wd batch size iterations token init optimizer lr scheduler
SD1.5 1e-3 1e-4 16 26,000 $\varnothing$ AdamW CosineAnnealingWarmRestarts
SDXL 1e-3 1e-4 16 30,000 $N(0,0.02)$ AdamW CosineAnnealingWarmRestarts
SD3 1e-3 1e-4 16 30,000 $N(0,0.02)$ AdamW CosineAnnealingWarmRestarts
Table 3: The implementation details for training.
Appendix C Prompt for Editing dataset
To generate source/target text prompt from images, we leverage LLaVA [19] [^1] We use checkpoint from 4bit/llava-v1.5-13b-3GB and Llama [8] [^2] We use Llama-3.1-8B-Instruct model. sequentially. Specifically, we extract source text description from 800 training images of DIV2K dataset by giving each image and following text prompt to LLaVA.
”Describe the object and background in the image”
Then, we generate the target text prompt that has only a single different concept compared with the source text prompt using Llama with the following instruction.
”You are an AI assistant for generating paired text prompts for real image editing tasks. Your goal is to modify a given text description by replacing an object with other while strictly following these rules:
  • 1) Modify only one object (i.e., a single meaningful concept such as an object). It could be small object.
  • 2) The replacement must be significantly different from the original concept but contextually appropriate. Avoid unrealistic substitutions (e.g., changing ”rabbit on grass” to ”rocket on grass”).
  • 3) Ensure diversity in word choices across different modifications.
  • 4) Preserve all other words exactly as they are. Do not change sentence structure, introduce new elements, or modify additional details.
  • 5) Do not provide any additional words—output only the modified text description.
  • 6) Do not change or add colors. Specifically, when modifying a building, change only the appearance, not the type of building (e.g., do not change ”building” to ”church” or ”lighthouse”).
  • 7) Modify only one feature at a time. If changing an object (e.g., ”starfish” to ”sea turtle”), do not alter its color, shape, or other attributes.
Example:
Input: The image features a close-up of a brown dog with a blue nose. The dog is standing in a grassy field, and the background is blurred, creating a focus on the dog’s face. The dog’s ears are perked up, and its eyes are open, giving it a curious and attentive expression. The dog’s fur is brown, and the grass in the background is green, creating a natural and vibrant scene.
Output: The image features a close-up of a brown fox with a blue nose. The fox is standing in a grassy field, and the background is blurred, creating a focus on the fox’s face. The fox’s ears are perked up, and its eyes are open, giving it a curious and attentive expression. The fox’s fur is brown, and the grass in the background is green, creating a natural and vibrant scene.”
Appendix D Additional Results on T2I Generation
The overall quantitative results are shown in Tab. 4. The additional qualitative results on SD1.5 and SDXL are shown in Fig. 8 and Fig. 9
Human Preference Text Alignment Image Quality
[…]
SD1.5 30 7.0 0.177 21.478 0.264 0.250 24.598 0.438
SD1.5 +Ours 30 7.0 0.273 21.541 0.270 0.252 25.358 0.437
SDXL 30 7.0 0.750 22.389 0.267 0.273 24.697 0.420
SDXL +Ours 30 7.0 0.852 22.620 0.268 0.283 26.040 0.423
SD3 28 4.0 0.942 22.549 0.263 0.280 31.595 0.424
SD3 +Ours 28 4.0 1.085 22.550 0.269 0.289 36.214 0.428
Table 4: Quantitative results of text to image generation on SD1.5, SDXL, and SD3.
Figure 8: The qualitative results of text-to-image generation comparing SD1.5 and SD1.5 with proposed method. The given text is from COCO dataset.
Figure 9: The qualitative results of text-to-image generation comparing SDXL and SDXL with proposed method. The given text is from COCO dataset.
Appendix E Additional Results on Image Editing
The overall quantitative results of DIV2K and Cat2Dog dataset on different target CFG scale are shown in Tab. 5 and Tab. 6, respectively. Additionally, qualitative results on different target CFG scale are shown in Tab. 5. The additional qualitative results across different number of editing steps are shown in Fig. 10
Figure 10: The additional qualitative results on editing using FlowEdit. The three different results adopting soft tokens (edited1-3) are vary on the number of editing steps, $27,29,30$ out of $50$ total timesteps, respectively.
Human Preference Text Alignment Image Quality
[…]
FlowEdit 33 13.5 0.380 21.749 0.261 0.255 52.038 0.154
FlowEdit 33 16.5 0.466 21.824 0.263 0.259 55.371 0.180
FlowEdit 33 19.5 0.510 21.869 0.265 0.261 58.069 0.200
FlowEdit +Ours 30 9 0.466 21.884 0.263 0.260 52.633 0.149
FlowEdit +Ours 30 11 0.564 21.985 0.268 0.265 56.900 0.178
FlowEdit +Ours 30 13 0.627 22.050 0.272 0.270 59.710 0.201
Table 5: Quantitative results of image editing regarding image quality and target text alignment of generated images with various target CFG scales on DIV2K dataset.
Human Preference Text Alignment Image Quality
[…]
FlowEdit 33 13.5 0.937 20.726 0.225 0.266 197.1 0.199
FlowEdit 33 16.5 0.908 20.641 0.229 0.272 201.05 0.217
FlowEdit 33 19.5 0.882 20.572 0.232 0.276 202.92 0.234
FlowEdit +Ours 30 7 1.111 21.143 0.226 0.269 198.71 0.173
FlowEdit +Ours 30 9 1.144 21.221 0.234 0.283 208.18 0.204
FlowEdit +Ours 30 11 1.144 21.224 0.239 0.290 214.72 0.230
FlowEdit +Ours 30 13 1.135 21.200 0.241 0.293 219.35 0.252
Table 6: Quantitative results of image editing regarding image quality and target text alignment of generated images with various target CFG scales on Cat2Dog dataset.
Appendix F Additional Results on Ablation Study
The overall quantitative results on different number of layers and number of tokens are shown in Tab. 7. Additionally, qualitative results are shown in Fig. 11 and Fig. 12, respectively. The qualitative results show that the text prompts are emphasized as more layers adopt soft tokens(Fig. 11). The same image of the ablation on the number of tokens are shown in Fig. 12.
Figure 11: Qualitative results of the ablation study on the number of layers adopting soft tokens.
Figure 12: Qualitative results of the ablation study on the number of tokens adopting soft tokens.
Human Preference Text Alignment Image Quality
[…]
8 2 0.993 22.556 0.263 0.255 72.253 0.428
8 4 1.009 22.464 0.266 0.255 72.308 0.424
8 6 0.984 22.403 0.267 0.255 73.840 0.427
8 7 0.944 22.197 0.269 0.255 74.546 0.425
1 5 1.054 22.492 0.272 0.283 58.430 0.426
4 5 1.063 22.493 0.269 0.287 60.766 0.429
8 5 1.056 22.516 0.268 0.288 62.644 0.431
16 5 0.801 22.095 0.265 0.278 60.743 0.431
32 5 0.675 21.876 0.257 0.275 61.623 0.426
Table 7: Quantitative results of image generation ablation study on the number of soft tokens and the number of layers.